OpenStudy (exceptionlearner):

if 1.01g of hydrogen(H) combines with 16.0g of oxygen (O), how many grams of hydrogen will be needed to combines with the same amount of oxygen to form H2O?

OpenStudy (rushwr):

What do u think the answer is ?

OpenStudy (exceptionlearner):

17.01 @Rushwr

OpenStudy (rushwr):

Show me ur work ?

OpenStudy (exceptionlearner):

16.0+1.01=17.01

OpenStudy (rushwr):

okai I will show u the work. We need one equation here n= m/M n - no. of moles m= mass M= molar mass S in the first part they say 1.01g of H combines with 16g of O no we will find the moles in each. ( this is not necessary since u seem to be confused only ) no. of moles of H = 1.01 /1.01 = 1 mole no. of moles of O = 16/16= 1 mole ( molar mass of O and H u have to know ) TO form H2O you need one mole of O and 2 moles of H. We know the moles of O present 16g (we found it earlier , 1mole) (moles of H = twice the no. of moles of O ) So now use the same equation for H We know the moles , molar mass and we have to find the mass If n= m/M then m = n * M right? So m = 2 * 1.01 = 2.02g

OpenStudy (exceptionlearner):

Oh...thank so much @Rushwr

OpenStudy (rushwr):

U got it right ?

OpenStudy (exceptionlearner):

Yes

OpenStudy (rushwr):

okai then have a nice day ^_^

OpenStudy (exceptionlearner):

You too @Rushwr